# When is the canonical conductor minimal?

### v1

@@ 3. Numerical semigroups with minimal canonical conductor
$\mathrm{QF}(H):=\{x\in\mathbb{N}\colon x+h\in H\text{ for all }h\in H\text{ with }h\neq 0\}$
and call an element of this set a quasi-Frobenius number. The Frobenius number $f$ is the largest quasi-Frobenius number.
@@
We can also see the type of $R=k[[H]]$ from the semigroup data. It is given by the size of $\mathrm{QF}(H)$. More precisely, if $f$ is the Frobenius number of $H$, then the $R$-module
$\omega=\sum_{x\in\mathrm{QF}(H)}Rt^{f-x}$
is a canonical module which is fractional ideal sitting between $R$ and $k[[t]]=\overline{R}$.
@@
Let $H$ be a numerical semigroup with Frobenius number $f$. Then, the ring $R=k[[H]]$ has minimal canonical conductor if and only if f1f-1 is a quasi-Frobenius number for $H$.
@@
If f1f-1 is a quasi-Frobenius number for $H$, then there exists a canonical fractional ideal $\omega=R+Rt+J$ where $J=R^{t_{1}}+\ldots+R^{t_{s}}$ for some $t_{1},\ldots,t_{s}$. Note that $\omega^{n}\subseteq k[[t]]=\overline{R}$ for any positive integer $n$. On the other hand, we have
@@
If f1f-1 is not a quasi-Frobenius number, we the canonical fractional ideal is of the form $\omega=R+R^{t_{1}}+\cdots+R^{t_{s}}$ for some $t_{1},\ldots,t_{s}\geq 2$. Therefore, we have $t\notin\omega^{n}$ for any $n\geq 1$. This finishes the proof. ∎
@@
Since the minimal number of generators equals the smallest generator, $k[[H]]$ is of minimal multiplicity. Therefore, quasi-Frobenius numbers of $H$ are given by $a_{2}-n,a_{3}-n,\ldots,a_{n}-n$. Hence, the Frobenius number is $a_{n}-n$ and $a_{n}-n-1$ is a Frobenius number if and only if $a_{n-1}=a_{n}-1$. ∎
@@
The inclusion $b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)$ is strict if and only if $n>3$. Hence this ring, while being nearly Gorenstein, is far from being almost Gorenstein for $n>3$.

### v2

@@ 5. Numerical semigroups with minimal canonical conductor
$\mathrm{PF}(H):=\{x\in\mathbb{Z}\setminus H\colon x+h\in H\text{ for all }h\in H\text{ with }h\neq 0\}$
and call an element of this set a pseudo-Frobenius number. The Frobenius number $f$ is the largest pseudo-Frobenius number.
@@
We can also see the type of $R=k[[H]]$ from the semigroup data. It is given by the size of $\mathrm{PF}(H)$. More precisely, if $f$ is the Frobenius number of $H$, then the $R$-module
$\omega=\sum_{x\in\mathrm{PF}(H)}Rt^{f-x}$
is a canonical module which is fractional ideal sitting between $R$ and $k[[t]]=\overline{R}$ [JÄG77, Korollar 5].
@@
Let $H$ be a numerical semigroup with Frobenius number $f$. Then, the ring $R=k[[H]]$ has minimal canonical conductor if and only if f1f-1 is a pseudo-Frobenius number for $H$.
@@
If f1f-1 is a pseudo-Frobenius number for $H$, then there exists a canonical fractional ideal $\omega=R+Rt+J$ where $J=R^{t_{1}}+\ldots+R^{t_{s}}$ for some $t_{1},\ldots,t_{s}$. Note that $\omega^{n}\subseteq k[[t]]=\overline{R}$ for any positive integer $n$. On the other hand, we have
@@
If f1f-1 is not a pseudo-Frobenius number, we the canonical fractional ideal is of the form $\omega=R+R^{t_{1}}+\cdots+R^{t_{s}}$ for some $t_{1},\ldots,t_{s}\geq 2$. Therefore, we have $t\notin\omega^{n}$ for any $n\geq 1$. This finishes the proof. ∎
@@
Since the minimal number of generators equals the smallest generator, $k[[H]]$ is of minimal multiplicity. Therefore, pseudo-Frobenius numbers of $H$ are given by $a_{2}-n,a_{3}-n,\ldots,a_{n}-n$. Hence, the Frobenius number is $a_{n}-n$ and $a_{n}-n-1$ is a Frobenius number if and only if $a_{n-1}=a_{n}-1$. ∎
@@
Therefore, $R$ has minimal canonical conductor. The inclusion $b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)$ is strict if and only if $n>3$. Hence this ring, while being nearly Gorenstein, is far from being almost Gorenstein for $n>3$.